# Action-Driven Flows for Causal Variational Principles

### v1

@@ 5. Further examples
We now illustrate the previous abstract results in a few examples. We choose $\mathscr{F}=\mathbb{R}^{2}$, $x_{0}\in\mathscr{F}$ and let $\rho_{0}:=\delta_{x_{0}}$ be the Dirac measure at $x_{0}$. Given a bounded continuous function $V\in\operatorname{C}^{0}(\mathscr{F},\mathbb{R})\cap L^{\infty}(\mathscr{F},\mathbb{R})$, we define the Lagrangian by
@@
We observe that, for a Dirac measure centered at some $x\in\mathscr{F}$, the the penalized action simplifies to
$\mathcal{S}^{h,\xi}(\delta_{x})=V(x)+\frac{1}{2h}|x-x_{0}|^{2}+\xi\>|x-x_{0}|\>,$ (5.2)
Since $V$ is bounded and continuous and $c>0$, this function is minimal at some $x_{1}\in\mathscr{F}$. Next, let $\rho\in\mathfrak{M}_{1}(\mathscr{F})$ be an arbitrary measure. Using that
@@
Now (5.3) is bounded from below by $\mathcal{S}^{h,\xi}(\delta_{x_{1}})$. Moreover, (5.4) is obviously non-negative, and it is zero if and only if $\rho$ is a Dirac measure. Finally (5.5) and (5.6) are non-negative in view of Hölder’s inequality for normalized measures (here we make essential use of the fact that $p\geq 2$). We conclude that every minimizing measure is a Dirac measure. ∎
@@ 6.1. Causal fermion systems and the reduced causal action principle
(causal fermion systems of fixed trace) Given a finite-dimensional Hilbert space $\mathscr{H}$ with scalar product $\langle.|.\rangle_{\mathscr{H}}$ and a parameter $n\in\mathbb{N}$ (the “spin dimension”), we we let $\mathscr{F}\subset\operatorname{L}(\mathscr{H})$ be the set of all symmetric linear operators $x$ on $\mathscr{H}$ with trace one,
@@
causal action: ${\mathcal{S}}(\rho)$ =×$=\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\>\operatorname{d}\!\rho(x)\,\operatorname{d}\!\rho(y)\>.$ (6.3)
@@
volume $\text{\em{volume constraint:}}\qquad\rho(\mathscr{F})=1\>,$
@@ 6.2. Moment measures and existence theory
For a given measure $\rho$ on $\mathscr{F}$, we define the measurable sets of ${\mathscr{K}}$ by the requirement that the sets $\mathbb{R}^{+}\Omega=\{\lambda p\>|\>\lambda\in\mathbb{R}^{+},p\in\Omega\}$ and $\mathbb{R}^{-}\Omega$ should be $\rho$-measurable in $\mathscr{F}$. We introduce the measures $\mathfrak{m}^{(0)}$, $\mathfrak{m}^{(1)}_{\pm}$ and $\mathfrak{m}^{(2)}$ by
@@
The measures $\mathfrak{m}^{(l)}$ and $\mathfrak{m}^{(l)}_{\pm}$ are referred to as the $l^{\text{th}}$ moment measure.
The main point is that the causal action as well as the constraints can be expressed purely in terms of the moment measures. Indeed, the volume constraint $\rho(\mathscr{F})=1$ and the trace constraints can be expressed as
@@
Here we made essential use of the fact that the trace is homogeneous of degree one and that the $\kappa$-Lagrangian in both arguments is homogeneous of degree two.
@@
with converge in the $\operatorname{C}^{0}({\mathscr{K}})^{*}$-topology, where $\mathfrak{m}^{(0)}\in\mathfrak{M}_{1}({\mathscr{K}})$ is a normalized Borel measure and $\mathfrak{m}^{(1)},\mathfrak{m}^{(2)}\in\mathfrak{M}({\mathscr{K}})$ are Borel measures. As shown in [6, Lemma 2.12] (for more details see also [10, Chapter 12]), we know that there is a parameter $\varepsilon$ (which depends only on the spin dimension $n$ and the dimension of the Hilbert space $f$) such that for any measurable set $\Omega\subset{\mathscr{K}}$ the following inequalities hold,
$\left|\mathfrak{m}^{(1)}(\Omega)\right|^{2}$ $\leqslant\mathfrak{m}(\Omega)\>\mathfrak{m}^{(2)}(\Omega)$ (6.6)
@@
The estimate (6.6) implies that $\mathfrak{m}^{(1)}$ is absolutely continuous with respect to $\mathfrak{m}$. Therefore, it has the Radon-Nikodym representation
@@ 6.3. Minimizing movements for the causal action principle
In view of the constructions of the previous section, it seems preferable to work with the moment measures. Moreover, the proof of Theorem 6.3 shows that, for constructing minimizers, it is no loss of generality to consider measures of the form
@@
$F:{\mathscr{K}}\rightarrow\mathscr{F}\>,\qquad x\mapsto f(x)\,x\qquad\text{with}\qquad f\in L^{2}({\mathscr{K}},\mathbb{R}^{+}_{0})$
(here and in what follows, we shall work only with the zeroth moment measure, which for simplicity will be denoted by $\mathfrak{m}$). According to (6.4), the volume and trace constraints are implemented by demanding that
@@
Since the $\kappa$-Lagrangian is non-negative, the penalized action is bounded below and thus $m:=\inf\mathcal{S}^{h,\xi}$ exists in $[0,\infty)$. We choose a minimizing sequence $(\mathfrak{m}_{\ell},f_{\ell})$ for $\mathcal{S}^{h,\xi}$, so that $m=\lim_{\ell\to\infty}\mathcal{S}^{h,\xi}(\mathfrak{m}_{\ell},f_{\ell})$. Due to the penalization, the sequences of measures $\mathfrak{m}_{\ell}$ and $|f_{\ell}^{q}|\,\mathfrak{m}_{\ell}$ are bounded. Therefore, the Banach-Alaoglu theorem provides us with a non-relabeled subsequence such that
@@
Now Theorem 4.6 and Proposition 4.8 extend in a straightforward way. The only additional ingredient to keep in mind is that the causal Lagrangian is indeed Hölder continuous with Hölder exponent $\alpha=1/(2n+1)$ (see [14, Theorems 5.1 and 5.3]), so that we can use the estimate (4.4).
@@
$\leqslant\frac{1}{\xi}\>\big{(}s_{2}-s_{1}\big{)}\qquad\text{for all~{}${\mathcal{S}}^{\xi}_{\min}<s_{1}<s_{2}\leqslant{\mathcal{S}}(\varrho_{0})$}\>.$
@@
Moreover, in analogy to Proposition 4.8, we have the following result.
Proposition 6.7.
$\leqslant\frac{1}{\xi}\>\big{(}s_{2}-s_{1}\big{)}\qquad\text{for all~{}${\mathcal{S}}^{\xi}_{\min}<s_{1}<s_{2}\leqslant{\mathcal{S}}(\varrho_{0})$}\>.$
Moreover, the limit $(\tilde{\mathfrak{m}}^{\xi}(s_{1}),\tilde{f})(\mathcal{S}_{\min}^{\xi}):=\mathrm{w}^{*}\text{-}\lim_{s\searrow{\mathcal{S}}_{\min}^{\xi}}\big{(}\tilde{\mathfrak{m}}^{\xi}(s),\tilde{f}^{\xi})(s)\big{)}$ exists in the sense of weak*-convergence of measures.
@@ 6.4. Limiting measures and Euler-Lagrange equations
In the case $\xi>0$, for any $q>0$ the curve $(\tilde{\mathfrak{m}}^{\xi}(s),\tilde{f}^{\xi}(s))$ converges with respect to the distance function (6.14). In the case of penalization by the Wasserstein distance $W_{p}$ (i.e. in Case 2. in (4.2)), the limiting measure
@@
is minimal on the support of $\rho$,
$\ell_{\xi}|_{N}\equiv\inf_{\mathscr{F}}\ell_{\xi}\>.$
@@
7. Application: A flow in the infinite-dimensional case
In order to exemplify possible applications of the constructed flows, we will now show how the Lipschitz continuous flow constructed in Proposition 6.7 can be used in order to construct a corresponding flow in the infinite-dimensional setting. The general idea is to append the flows in finite-dimensional subspaces of the Hilbert space for increasing dimension. For the detailed construction, we assume that the Hilbert space $\mathscr{H}$ in Definition 6.1 is separable but $\dim\mathscr{H}=\infty$. We consider a filtration by finite-dimensional subspaces, i.e.
$\mathscr{H}_{1}\subset\mathscr{H}_{2}\subset\cdots\subset\mathscr{H}\quad\text{with}\quad\dim\mathscr{H}_{p}=p\qquad\text{and}\qquad\mathscr{H}=\overline{\bigcup_{p=1}^{\infty}\mathscr{H}_{p}\!\!\!}^{\langle.|.\rangle_{\mathscr{H}}}\>.$
Extending the operators by zero, we obtain corresponding inclusions $\mathscr{F}_{1}\subset\mathscr{F}_{2}\subset...\subset\mathscr{F}$ with
@@ 7. Application: A flow in the infinite-dimensional case
Using the above embeddings, we can consider this limiting measure as being in ${\mathcal{P}}({\mathscr{K}})$. Taking this measure as the new starting point, we consider the reparametrized flow from Proposition 6.7 in $\mathscr{F}_{2}$. Proceeding in this way inductively, we obtain a Lipschitz continuous curve in $\mathscr{F}$. The flow is strictly decreasing along the curve.
We finally mention that the above method can be refined in various ways. One extension which seems useful is not to choose $\xi>0$ as a constant, but to consider instead a monotone decreasing sequence $(\xi_{p})_{p\in\mathbb{N}}$ which converges to zero as the dimension $p$ of the Hilbert space tends to infinity.
Acknowledgments: F.G. would like to thank the Hector Foundation for support.

### v2

@@ 5. Further examples
We now illustrate the previous abstract results in a few examples. We choose $\mathscr{F}=\overline{B_{1}(0)}\subset\mathbb{R}^{2}$ as a closed unit ball in two dimensions. Moreover, we choose $x_{0}\in\mathscr{F}$ and let $\rho_{0}:=\delta_{x_{0}}$ be the Dirac measure at $x_{0}$. Given a bounded continuous function $V\in\operatorname{C}^{0}(\mathscr{F},\mathbb{R})\cap L^{\infty}(\mathscr{F},\mathbb{R})$, we define the Lagrangian by
@@
We observe that, for a Dirac measure centered at some $x\in\mathscr{F}$, the penalized action simplifies to
$\mathcal{S}^{h,\xi}(\delta_{x})=V(x)+\frac{1}{2h}|x-x_{0}|^{2}+\xi\>|x-x_{0}|\>.$ (5.2)
Since $V$ is bounded and continuous, this function is minimal at some $x_{1}\in\mathscr{F}$. Next, let $\rho\in\mathfrak{M}_{1}(\mathscr{F})$ be an arbitrary measure. Using that
@@
Now (5.3) is bounded from below by $\mathcal{S}^{h,\xi}(\delta_{x_{1}})$ (recall that $x_{1}$ was defined as the minimizer of the integrand of (5.3)). Moreover, (5.4) is obviously non-negative, and it is zero if and only if $\rho$ is a Dirac measure. Finally, the summands in (5.5) and (5.6) are non-negative in view of Hölder’s inequality for normalized measures (here we make essential use of the fact that $p\geq 2$). We conclude that every minimizing measure is a Dirac measure. ∎
@@ 6.1. Causal fermion systems and the reduced causal action principle
(causal fermion systems of fixed trace) Given a finite-dimensional Hilbert space $\mathscr{H}$ with scalar product $\langle.|.\rangle_{\mathscr{H}}$ and a parameter $n\in\mathbb{N}$ (the “spin dimension”), we let $\mathscr{F}\subset\operatorname{L}(\mathscr{H})$ be the set of all symmetric linear operators $x$ on $\mathscr{H}$ with trace one,
@@
causal action: ${\mathcal{S}}(\rho)$ $=\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\>\operatorname{d}\!\rho(x)\,\operatorname{d}\!\rho(y)\>.$ (6.3)
@@
$\text{{volume constraint:}}\qquad\rho(\mathscr{F})=1\>,$
@@ 6.2. Moment measures and existence theory
For a given measure $\rho$ on $\mathscr{F}$, we define the measurable sets $\Omega\subset{\mathscr{K}}$ by the requirement that the sets $\mathbb{R}^{+}\Omega=\{\lambda p\>|\>\lambda\in\mathbb{R}^{+},p\in\Omega\}$ and $\mathbb{R}^{-}\Omega$ should be $\rho$-measurable in $\mathscr{F}$. We introduce the measures $\mathfrak{m}^{(0)}$, $\mathfrak{m}^{(1)}_{\pm}$ and $\mathfrak{m}^{(2)}$ by
@@
The measures $\mathfrak{m}^{(l)}$ and $\mathfrak{m}^{(l)}_{\pm}$ are referred to as the $l^{\text{th}}$ moment measures.
The main point is that the causal action as well as the constraints can be expressed purely in terms of the moment measures. Indeed, as shown in [6, Section 2.3] (for more details see also [10, Section 12.6]), the volume constraint $\rho(\mathscr{F})=1$ and the trace constraints can be expressed as
@@
Here we make essential use of the fact that the trace is homogeneous of degree one and that the $\kappa$-Lagrangian in both arguments is homogeneous of degree two.
@@
with convergence in the $\operatorname{C}^{0}({\mathscr{K}})^{*}$-topology, where $\mathfrak{m}^{(0)}\in\mathfrak{M}_{1}({\mathscr{K}})$ is a normalized Borel measure and $\mathfrak{m}^{(1)}_{\pm},\mathfrak{m}^{(2)}\in\mathfrak{M}({\mathscr{K}})$ are Borel measures. As shown in [6, Lemma 2.12] (for more details see also [10, Chapter 12]), we know that there is a parameter $\varepsilon$ (which depends only on the spin dimension $n$ and the dimension of the Hilbert space $f$) such that for any measurable set $\Omega\subset{\mathscr{K}}$ the following inequalities hold,
$\mathfrak{m}^{(1)}_{\pm}(\Omega)^{2}$ $\leqslant\mathfrak{m}(\Omega)\>\mathfrak{m}^{(2)}(\Omega)$ (6.6)
@@
These inequalities show that the measures $\mathfrak{m}^{(2)}$ and $\mathfrak{m}^{(1)}_{\pm}$ are bounded. Therefore, we can introduce the signed measure $\mathfrak{m}^{(1)}$ by $\mathfrak{m}^{(1)}:=\mathfrak{m}^{(1)}_{+}-\mathfrak{m}^{(1)}_{-}$. The estimate (6.6) implies that this signed measure is absolutely continuous with respect to $\mathfrak{m}$. Therefore, it has the Radon-Nikodym representation
@@ 6.3. Minimizing movements for the causal action principle
In view of the constructions of the previous section, it seems preferable to work with the moment measures. For notational simplicity, we denote the zeroth moment measure by $\mathfrak{m}$. Then the proof of Theorem 6.3 shows that, for constructing minimizers, it is no loss of generality to consider measures of the form
@@
[…]
According to (6.4), the volume and trace constraints are implemented by demanding that
@@
Since the $\kappa$-Lagrangian is non-negative, the penalized action is bounded below and thus $m:=\inf\mathcal{S}^{h,\xi}$ exists in $0,\infty)$. We choose a minimizing sequence $(\mathfrak{m}_{\ell},f_{\ell})$ for $\mathcal{S}^{h,\xi}$, so that $m=\lim_{\ell\to\infty}\mathcal{S}^{h,\xi}(\mathfrak{m}_{\ell},f_{\ell})$. Due to the penalization, the sequences of measures $\mathfrak{m}_{\ell}$ and $|f_{\ell}^{q}|\,\mathfrak{m}_{\ell}$ are bounded. Therefore, the Banach-Alaoglu theorem provides us with a non-relabeled subsequence such that
@@
Now Propositions 4.6 and 4.8 extend in a straightforward way. The only additional ingredient to keep in mind is that the causal Lagrangian is indeed Hölder continuous with Hölder exponent $\alpha=1/(2n+1)$ (see [14, Theorems 5.1 and 5.3]), so that we can use the estimate (4.4).
@@
$\leqslant\frac{1}{\xi}\>\big(s_{2}-s_{1}\big)\qquad\text{for all~${\mathcal{S}}^{\xi}_{\min}\leqslant s_{1}<s_{2}\leqslant{\mathcal{S}}(\varrho_{0})$}\>.$
@@ 6.4. Limiting measures and Euler-Lagrange equations
In the case $\xi>0$, for any $q>0$ the curve $(\tilde{\mathfrak{m}}^{\xi}(s),\tilde{f}^{\xi}(s))$ converges with respect to the distance function (6.14) as $s\searrow{\mathcal{S}}^{\xi}_{\min}$. In the case of penalization by the Wasserstein distance $W_{p}$ (i.e. in Case 2. in (4.2)), the limiting measure
@@
is minimal on the support of $\rho$, i.e.
$\ell_{\xi}|_{N}\equiv\inf_{\mathscr{F}}\ell_{\xi}$
with $N:=\operatorname{supp}\rho$ and $\rho$ defined similar to (6.9) by $\rho:=\tilde{F}_{*}\tilde{m}^{\xi}$ and $\tilde{F}(x):=\tilde{f}(x)\,x$.
@@
7. Application and outlook: A flow in the infinite-dimensional case
In order to exemplify possible applications of the constructed flows, we will now show how the Lipschitz continuous flow constructed in Proposition 6.6 can be used in order to construct a corresponding flow in the infinite-dimensional setting. The general idea is to append the flows in finite-dimensional subspaces of the Hilbert space for increasing dimension.
For the detailed construction, we assume that the Hilbert space $\mathscr{H}$ in Definition 6.1 is separable but $\dim\mathscr{H}=\infty$. We consider a filtration by finite-dimensional subspaces, i.e.
$\mathscr{H}_{1}\subset\mathscr{H}_{2}\subset\cdots\subset\mathscr{H}\quad\text{with}\quad\dim\mathscr{H}_{p}=p\qquad\text{and}\qquad\mathscr{H}=\overline{\bigcup_{p=1}^{\infty}\mathscr{H}_{p}\!\!\!}^{\langle.|.\rangle_{\mathscr{H}}}\>.$ (7.1)
Extending the operators by zero, we obtain corresponding inclusions $\mathscr{F}_{1}\subset\mathscr{F}_{2}\subset$$\subset\mathscr{F}$ with
@@ 7. Application and outlook: A flow in the infinite-dimensional case
Using the above embeddings, we can consider this limiting measure as being in ${\mathcal{P}}({\mathscr{K}})$. Taking this measure as the new starting point, we consider the reparametrized flow from Proposition 6.6 in $\mathscr{F}_{2}$. Proceeding in this way inductively, we obtain a Lipschitz continuous curve in $\mathscr{F}$. The action is strictly decreasing along the flow curve.
We note that the above method can be refined in various ways. One extension which seems useful is not to choose $\xi>0$ as a constant, but to consider instead a monotone decreasing sequence $(\xi_{p})_{p\in\mathbb{N}}$ which converges to zero as the dimension $p$ of the Hilbert space tends to infinity. Similarly one can also adjust the parameter $\kappa$ in (6.2) when increasing the dimension. The detailed construction remains to be worked out.
We finally remark that this procedure is inspired by and bears some resemblance with renormalization flow techniques used in quantum field theory, In order to explain the connection, we note that ultraviolet regularizations are often realized by a cutoff in momentum space which (at least for systems in finite spatial volume) corresponds to restricting attention to finite-dimensional subspaces of the underlying Hilbert space. Removing the cutoff corresponds to the limit when the dimensions of the subspaces tends to infinity. In the renormalization program, one studies this limit while carefully adjusting the masses and coupling constants in the physical action. Our analysis is similar because we study minimizers of the causal action for a filtration (7.1) while adjusting the parameters $\xi$ and $\kappa$.
Acknowledgments: We would like to thank the referees for the careful reading and many useful suggestions. F.G. would like to thank the Hector Foundation for support.